OpenStudy (anonymous):

Calculate the hydroxide ion concentration of a solution with pH = 10.25. Show all calculations leading to an answer.

OpenStudy (studygurl14):

-log[H+] = pH

OpenStudy (studygurl14):

You know that pH = 10.25. Plug into the equation -log[H+] = 10.25 Now solve for H+ log[H+] = -10.25 You'll need to do antilog on your calculator (which is \(10^x\)) H+ = \(10^{-10.25}\)

OpenStudy (studygurl14):

Find \(\large 10^{-10.25}\) Use a calculator.

OpenStudy (anonymous):

The Kw of water varies with temperature. Calculate the pH of water at 46⁰C with a Kw = 3.219 x 10-14. Show all calculations leading to an answer. @StudyGurl14

OpenStudy (anonymous):

[H+] * [OH-] = 3.219 x 10-14 [H+] = 1.79 * 10^-7 pH = - log[H+] = 6.75 is that right?

OpenStudy (studygurl14):

i think so...i'm not sure though...

OpenStudy (anonymous):

If 6.40 kcal of heat is added to 1.00 kg of ice at 0⁰C, how much liquid water at 0⁰C is produced? Show all calculations leading to an answer. @StudyGurl14

OpenStudy (anonymous):

Use the latent heat of fusion for ice and \[\color{green}{E = ml_f}\] Solve for m:\[\color{green}{m = \dfrac{E}{l_f} = \dfrac{6.40~\text{kcal}}{l_f}}\]

OpenStudy (anonymous):

Look up the value of the latent heat of fusion for water. Put it in the formula for \[\color{green}{l_f}\]

OpenStudy (anonymous):

Did you look up the value of the latent heat of fusion as I suggested?

OpenStudy (anonymous):

336000 @ghuczek

OpenStudy (anonymous):

What are the units?

OpenStudy (anonymous):

336000 J per kg

OpenStudy (anonymous):

You will need to do a unit conversion. Convert 6.40 kcal to joules.

OpenStudy (anonymous):

Multiply by 4184. Then when you have done that, divide the answer you get by 336000 and it will give you the mass of the water, in kilograms ... and you are done!

OpenStudy (anonymous):

youre amazing! can you help with more? @ghuczek

OpenStudy (anonymous):

An aqueous solution freezes at -3.47⁰C. What is its boiling point? Show all calculations leading to an answer. @ghuczek

OpenStudy (anonymous):

I haven't got time right now, sorry. I have to go.